OpenStudy (anonymous):

(x-5)^2-((y+1)^2/4)=1 This is supposed to be a graph of a hyperbola and I'm supposed to find the vertices and foci. I'm really confused, please help!

OpenStudy (mertsj):

What is the center?

OpenStudy (anonymous):

I believe it would be (5,-1)

OpenStudy (mertsj):

That is correct. What is the value of a^2?

OpenStudy (anonymous):

Where did the a come from?

OpenStudy (mertsj):

Does this hyperbola open to the right and left or up and down?

OpenStudy (anonymous):

Up?

OpenStudy (mertsj):

How do you know? Was that a guess?

OpenStudy (anonymous):

Well, the x came first so its up or down. And its positive.

OpenStudy (mertsj):

Since x is positive, it opens right and left.

OpenStudy (anonymous):

Is that because its on the x axis?

OpenStudy (mertsj):

(x-h)^2/a^2-(y-k)^2/b^2=1

OpenStudy (mertsj):

It's because x cannot be 0

OpenStudy (mertsj):

The equation I posted is the equation of a hyperbola with center (h,k) that opens right and left. What is the value of a^2 in your problem?

OpenStudy (anonymous):

1?

OpenStudy (mertsj):

and if a^2 is 1 then a is also 1 and the vertices are 1 unit to the right and 1 unit to the left of the center. So find the vertices.

OpenStudy (anonymous):

(4,-1), (6,-1)

OpenStudy (mertsj):

yes. Very good

OpenStudy (mertsj):

Now for a hyperbola that opens right and left, the foci are c units to the right and c units to the left of the center.

OpenStudy (mertsj):

The relationship between a, b, and c is that a^2+b^2=c^2 We know that a^2=1. What is b^2?

OpenStudy (anonymous):

How do I find c?

OpenStudy (mertsj):

The relationship between a, b, and c is that a^2+b^2=c^2 We know that a^2=1. What is b^2?

OpenStudy (anonymous):

c^2-1

OpenStudy (mertsj):

What is b^2?

OpenStudy (anonymous):

How do I find b without knowing what c is?

OpenStudy (mertsj):

The equation of a hyperbola that opens to the right and the left and has center (h,k) is: (x-h)^2/a^2-(y-k)^2/b^2=1

OpenStudy (mertsj):

Compare that to your equation and find out what b^2 is

OpenStudy (anonymous):

so b^2=4 and b=2

OpenStudy (mertsj):

Yes. a^2=1 and b^2=4 Now find c knowing that a^2+b^2=c^2

OpenStudy (anonymous):

c^2=5 so c=

OpenStudy (mertsj):

c=sqrt5

OpenStudy (anonymous):

the square root of 5

OpenStudy (mertsj):

and the foci are sqrt5 units to the right and left of the center so they are???

OpenStudy (anonymous):

(5+sqrt5, -1) and (5-sqrt5, -1)

OpenStudy (mertsj):

Very good.